How much work needed to stretch an extra 2 cm?
l m RS CWT nel de -to -stre tan
Answer #1

Energy stored in spring = 0.5*K*x2

At 5 cm energy stored = 0.5*160*0.05*.0.05 = 0.2 joule

At 3 cm = 0.5*160*0.03*0.03 = 0.072 joule

Therefore work done = 0.2 - 0.072= 0.128 joule